An expression for the amplitude of a resultant standing wave on a vibrating string is 2Asinkx. Show that the maximum values of the amplitude of such wave at location x=(2n-1) times labda/2.


1
Expert's answer
2021-03-29T08:59:42-0400

The extrema of function "f(x) = 2A\sin (kx)" are located at the points:

"f'(x_{e}) = 0\\
2kA\cos (kx_e) = 0"

The solutions of the last equation is:


"x_e = \dfrac{\pi}{2k} + \dfrac{m\pi}{k},\space\space\space m = 0, \pm1, \pm2, ..."

Since the wave is standing one, the locations of maxima and minima are the same. Thus, we can write for the maxima points:


"x_m = \dfrac{\pi}{2k} + \dfrac{n\pi}{k},\space\space\space n = 0, \pm1, \pm2, ...\\
x_m =\dfrac{\pi(2n+1)}{2k}"

Since

"k = \dfrac{2\pi}{\lambda}"

obtain:


"x_m = \dfrac{(2n+1)\lambda }{4} = \left(2n+1\right)\dfrac{\lambda}{4}"

Answer. The actual expression for the maxima is "x_m = \left(2n+1\right)\dfrac{\lambda}{4}".